# Energetic Electrons Accelerated and Trapped in a Magnetic Bottle above a Solar Flare Arcade

### v1

@@ 2.2 EUV Plasma Downflows
In the ALT magnetic bottle region, the magnetic field lines associated with the downward reconnection outflows are expected to start from a nearly anti-parallel configuration in the reconnection current sheet to a cusp-like shape after they pass the Y point. Then, the outflows quickly slow down until they join the flare arcade and relax to a potential state with a loop-like shape. This process is clearly shown in the SDO/AIA 171 Å EUV time-series images. In Figures 7(b)–(e), to better present the dynamic features, we plot a series of running-ratio images, each showing the intensity ratio of the current image to the one from 36-s earlier. There, a downward moving feature at the central current sheet location is seen to transform from an initially linear shape to a cusp shape as it moves across a location at $y\approx 43$ Mm. Such a transformation is already suggestive of the presence of a Y point at this location.
@@
Furthermore, in Figure 7(g), we show a space-time diagram obtained at a vertical slice located at the central current sheet. Multiple downward-moving features can be distinguished as tracks that orient toward the lower-right direction. The example downflow event shown in panels (b)–(e) has an initial speed of $>$900 km s-1 and quickly slows down to only $\approx 200$ km s-1 after it passes the same location. Such a sudden slow-down motion of plasma downflows further supports the presence of the Y point. Coincidentally, the Y point identified with the EUV imaging observations alone matches almost exactly with the location of the ALT HXR source shown in Figure 3. It is also fully consistent with the location of the Y-point-hosting magnetic bottle structure inferred from the microwave spectral imaging analysis, which has a weaker magnetic field strength.
@@
We note that the plane-of-sky-projected speed of the observed plasma downflow above the Y point ($\sim$900 km s-1) seems slower than the inferred local Alfvén speed in the ALT magnetic bottle region ($v_{A}\approx 1,900$ km s-1 for $B\approx 150$ G, , and $n_{\rm th}\approx 3\times 10^{10}$ cm-3, . By combining line-of-sight (LOS) flow measurements made by Hinode/EIS and plane-of-sky measurements from SDO/AIA, French et al. (2024) reported that despite that the reconnection current sheet appears to be viewed perfectly edge-on, the non-zero LOS speeds in the downflow region suggest that the current sheet is tilted slightly away from the observer. Therefore, the physical flow speed should be greater than that measured in the plane of the sky, although the required projection correction is likely insignificant given the low LOS speed ($<$35 km s-1) measured by Hinode/EIS. Moreover, it has been shown in recent 3D modeling that the speed of the observed plasma outflows in EUV time-series images may be substantially slower than the intrinsic reconnection outflows (Shen et al., 2023). Lastly, as argued by Chen et al. (2020b), the low cadence of SDO/AIA (12 s) may simply render flows faster than $\sim$2,000 km s-1 undetectable within the flaring region because they would traverse a large distance of $>$50 Mm in two frames. For these reasons, we argue that super-magnetosonic outflows may still be present to drive a fast-mode termination shock in the magnetic bottle region to facilitate the acceleration of the energetic particles.
@@ 3 MHD, Particle, and Emission Modeling
Figure 8: Distribution of nonthermal electrons around the magnetic bottle region and the associated emissions. (a) and (b) Spatial distribution of $>$50 keV and $>$300 keV electrons, respectively. Note they are shown on a logarithmic scale. (c) Example model nonthermal electron spectra from regions marked in panels (a) and (b), with one located near the Y point (red), two in the loop legs (green and cyan), . (d) Contours of $>$50 keV electron distribution overlaid on the model $v_{y}$ map. Blue, green, and red contours are 0.1%, 1%, and 15% of the peak electron density, respectively. In particular, the innermost red contour encloses 51% of all the $>$50 keV electrons. (e) Synthetic SDO/AIA (background), 25–60 keV HXR (red contours), and multi-frequency microwave sources (filled color contours) as calculated from the combined MHD and particle model. (d) Multi-frequency observations identical to Figure 3(d), overlaid with magnetic field lines derived from the model.
The multi-wavelength analysis above suggests that the ALT magnetic bottle region is where the majority of the microwave- and X-ray-emitting nonthermal electrons are concentrated. To further elucidate the role of the magnetic bottle and the underlying physical processes, we combine a data-informed macroscopic MHD and particle model to simulate the acceleration and transport of energetic electrons in a realistic flare geometry. The resistive 2.5-D MHD simulation we perform here is similar to those used in our earlier works (Chen et al., 2015, 2019; Shen et al., 2018), which features a dynamic fast-mode termination shock in the ALT magnetic bottle region. Appropriate scaling is applied to the MHD model to adjust to the observed flare size and observational constraints of the plasma properties. To model the electron acceleration and transport processes in the macroscopic flare geometry, we adopt the method used in Kong et al. (2019) by injecting pseudo electrons into the MHD model and simulating their kinetic evolution by solving the Parker transport equation. In the model, the electrons are found to be primarily accelerated via the Fermi mechanism due to compression (Fermi, 1949; Parker, 1965) in the magnetic bottle region, where the downward plasma flows collide head-on with the newly reconnected field lines. The fast-mode termination shock further facilitates the acceleration thanks to the sharp jump of physical parameters across the shock surface (Kong et al., 2019). We refer interested readers to Appendix Section C and references therein for more technical details.
@@
Figure 9: Comparison between the observed and modeled microwave images and spectra. (a) Observed multi-frequency microwave images overlaid on SDO/AIA 131 Å EUV image (identical to Figure 8(f)). (b) Sample microwave spectra derived from four selected regions marked in (a). (c) and (d) are identical to (a) and (b), but instead showing the modeled multi-frequency microwave images and the corresponding sample spectra.
@@
By combining the resulting distribution of energetic electrons from the particle model $dn_{e}/d\varepsilon(x,y,\varepsilon)$ and plasma properties including magnetic field $B(x,y)$, thermal plasma density $n_{\rm th}(x,y)$, and temperature $T(x,y)$ from the MHD model, we generate synthetic SDO/AIA EUV, RHESSI HXR, and EOVSA microwave images at different energy/frequency bands. Appropriate emission mechanisms are assumed for each emission type: optically thin line emissions for EUV, thin-target bremsstrahlung for HXRs, and nonthermal gyrosynchrotron emission for microwaves. The corresponding instrument response is also considered (see Appendix Section D for details). Figure 8(e) shows a composite of the resulting synthetic SDO/AIA 131 Å EUV image (background), RHESSI 25–60 keV HXR image (red contours), and multi-frequency EOVSA microwave images (filled color contours). All of the images display a striking resemblance to the observations shown in Figure 8(f): The EUV 131 Å image displays a bright, closed arcade with a cusp-shaped top. A compact HXR source is located in the magnetic bottle region near the fast-mode termination shock above the flare arcade. Meanwhile, similar to the observations, the multi-frequency microwave source resembles an arcade-like shape at the outer rim of the flare arcade. Figure 5 shows a side-by-side comparison between the observed and modeled microwave images at the same frequencies. Figures 8(e) and (f) display another comparison with the microwave images shown as overlaying filled contours colored from red to blue for increasing frequencies. One can see the model images resemble the observed ones remarkably well. Moreover, as shown in Figure 9, not only does the modeled microwave source morphology achieve an excellent match with the observations, but the spatially resolved microwave spectra derived from different regions also yield a qualitative agreement, demonstrating the success of our approach. The microwave source’s appearance is very different from its HXR counterpart mainly because, unlike the X-ray bremsstrahlung, gyrosynchrotron radiation has a strong dependence on the local magnetic field.
@@
Intriguingly, unlike the HXR source that is highly localized at the ALT magnetic bottle, similar to the observations, the multi-frequency microwave source extends well beyond the region toward the directions of the current sheet and the two footpoints. Such a large extension suggests that, while the majority of the energetic electrons are confined within the ALT HXR source near the Y point—in the model, it contains $\sim$50of the total $>$50 keV electrons (see red contour in Figure 8(d))—a small fraction of these electrons manage to escape and spread beyond the magnetic bottle region (albeit “diluted” to a much smaller density), giving rise to the extended microwave source with an arcade-like shape.
@@
We note that in the observations, another strong looptop X-ray source is present, whose centroid is located at $y\approx 24$ Mm (Figure 8(f)). Spectral analysis results suggest that it is extremely hot (15–27 MK) and dense (up to $10^{12}$ cm-3). This source has been interpreted as the result of thermalization of previously accelerated nonthermal particles (Veronig & Brown, 2004), colliding chromospheric upflows (Reeves et al., 2007), or compression by slow shocks formed by sheared reconnection (Longcope & Guidoni, 2011). Reproducing this additional source in a model requires the inclusion of feedback between the thermal plasma and nonthermal particles, more accurate treatment of the chromospheric evaporation processes, and modeling of the reconnection processes in the third dimension, which is beyond the scope of our current work.
@@ 4 Discussion and Conclusion
We have demonstrated that our model, which includes both MHD and particle processes in a realistic flare geometry, can reproduce emission signatures that are well-matched to multi-wavelength observations. In the model, the energetic electrons are found to be primarily accelerated in the magnetic bottle region by the converging flows and facilitated by the fast-mode termination shock via the Fermi mechanism. The accelerated electrons are further trapped there due to pitch-angle scattering by turbulence. This is the first time when synthetic observables in both the thermal and nonthermal regimes are generated from a self-consistent, macroscopic numerical model to compare with microwave, EUV, and X-ray imaging and spectroscopy observations. Such a remarkable agreement between the modeled and observed emissions suggests that this model of electron acceleration and transport is a viable approach during this period of interest.
@@
Figure 10: Physical picture suggested by our combined observation-modeling results. Energetic electrons are accelerated via the Fermi mechanism and trapped in the above-the-looptop magnetic bottle region due to turbulence. The reconnected and relaxed magnetic loops filled with hot plasma are observed as a bright EUV flare arcade with a hot looptop HXR source. The highly concentrated energetic electrons near the Y point produce the above-the-looptop HXR source via bremsstrahlung. Meanwhile, a small fraction of energetic electrons escaping from the magnetic bottle region produces the observed multi-frequency microwave source with a large extension, which has an arcade-like shape bestriding above the EUV flare arcade.
@@
The escaped nonthermal electrons gyrate in the magnetic field and produce microwave sources that form an arcade-like shape bestriding above the bright EUV flare arcade. The absence of an HXR footpoint source in our case suggests that it is either partially occulted by the solar limb at the time or the flux of nonthermal electrons reaching the footpoints is insufficient to produce observable HXR footpoint emission in the presence of a bright looptop source. In other cases when HXR footpoint sources are co-observed, they are often found to be located at the outer edge of bright EUV flare arcades (e.g., Liu, 2013; Krucker & Battaglia, 2014). These observations nicely corroborate our scenario: they result from accelerated nonthermal electrons escaping from the ALT magnetic bottle region and reaching the footpoints along the freshly reconnected field lines (see, e.g., recent modeling results by Kong et al., 2022b).

### v2

@@ 2.2 EUV Plasma Downflows
In the ALT magnetic bottle region, the magnetic field lines associated with the downward reconnection outflows are expected to start from a nearly antiparallel configuration in the reconnection current sheet to a cusplike shape after they pass the Y point. Then, the outflows quickly slow down until they join the flare arcade and relax to a potential state with a looplike shape. This process is clearly shown in the SDO/AIA 171 Å EUV time-series images. In Figures 7(b)–(e), to better present the dynamic features, we plot a series of running-ratio images, each showing the intensity ratio of the current image to the one from 36 s earlier. There, a downward-moving feature at the central current sheet location is seen to transform from an initially linear shape to a cusp shape as it moves across a location at $y\approx 43$ Mm. Such a transformation is already suggestive of the presence of a Y point at this location.
@@
Furthermore, in Figure 7(g), we show a space-time diagram obtained at a vertical slice located at the central current sheet. Multiple downward-moving features can be distinguished as tracks that orient toward the lower right. The example downflow event shown in panels (b)–(e) has an initial speed of $>$900 km s-1 and quickly slows down to only $\approx 200$ km s-1 after it passes the same location. Such a sudden slowdown motion of plasma downflows further supports the presence of the Y point. Coincidentally, the Y point identified with the EUV imaging observations alone matches almost exactly with the location of the ALT HXR source shown in Figure 3. It is also fully consistent with the location of the Y-point-hosting magnetic bottle structure inferred from the microwave spectral imaging analysis, which has a weaker magnetic field strength.
@@
We note that the plane-of-sky-projected speed of the observed plasma downflow above the Y point ($\sim$900 km s-1) seems slower than the inferred local Alfvén speed in the ALT magnetic bottle region ($v_{\rm A}\approx 1,900$ km s-1 for $B\approx 150$ G, constrained by our microwave spectral imaging analysis, and $n_{\rm th}\approx 3\times 10^{10}$ cm-3, constrained by Hinode/EIS measurements discussed in Appendix C). By combining line-of-sight (LOS) flow measurements made by Hinode/EIS and plane-of-sky measurements from SDO/AIA, French et al. (2024) reported that despite that the reconnection current sheet appearing to be viewed perfectly edge-on, the nonzero LOS speeds in the downflow region suggest that the current sheet is tilted slightly away from the observer. Therefore, the physical flow speed should be greater than that measured in the plane of the sky, although the required projection correction is likely insignificant given the low LOS speed ($<$35 km s-1) measured by Hinode/EIS. Moreover, it has been shown in recent 3D modeling that the speed of the observed plasma outflows in EUV time-series images may be substantially slower than the intrinsic reconnection outflows (Shen et al., 2023). Lastly, as argued by Chen et al. (2020b), the low cadence of SDO/AIA (12 s) may simply render flows faster than $\sim$2000 km s-1 undetectable within the flaring region because they would traverse a large distance of $>$50 Mm in two frames. For these reasons, we argue that supermagnetosonic outflows may still be present to drive a fast-mode termination shock in the magnetic bottle region to facilitate the acceleration of the energetic particles.
@@ 3 MHD, Particle, and Emission Modeling
Figure 8: Distribution of nonthermal electrons around the magnetic bottle region and the associated emissions. (a) and (b) Spatial distribution of $>$50 keV and $>$300 keV electrons, respectively. Note that they are shown on a logarithmic scale. (c) Example model nonthermal electron spectra from four regions marked in panels (a) and (b), with one located near the Y point (red), two in the loop legs (green and cyan), and one in the current sheet region (blue). (d) Contours of $>$50 keV electron distribution overlaid on the model $v_{y}$ map. Blue, green, and red contours are 0.1%, 1%, and 15% of the peak electron density, respectively. In particular, the innermost red contour encloses 51% of all the $>$50 keV electrons. (e) Synthetic SDO/AIA (background), 25–60 keV HXR (red contours), and multifrequency microwave sources (filled color contours) as calculated from the combined MHD and particle model. (f) Multifrequency observations identical to Figure 3(d), overlaid with magnetic field lines derived from the model.
The multiwavelength analysis above suggests that the ALT magnetic bottle region is where the majority of the microwave- and X-ray-emitting nonthermal electrons are concentrated. To further elucidate the role of the magnetic bottle and the underlying physical processes, we combine a data-informed macroscopic MHD and particle model to simulate the acceleration and transport of energetic electrons in a realistic flare geometry. The resistive 2.5D MHD simulation we perform here is similar to those used in our earlier works (Chen et al., 2015, 2019; Shen et al., 2018), which features a dynamic fast-mode termination shock in the ALT magnetic bottle region. Appropriate scaling is applied to the MHD model to adjust to the observed flare size and observational constraints of the plasma properties. To model the electron acceleration and transport processes in the macroscopic flare geometry, we adopt the method used in Kong et al. (2019) by injecting pseudoelectrons into the MHD model and simulating their kinetic evolution by solving the Parker transport equation. In the model, the electrons are found to be primarily accelerated via the Fermi mechanism due to compression (Fermi, 1949; Parker, 1965) in the magnetic bottle region, where the downward plasma flows collide head-on with the newly reconnected field lines. The fast-mode termination shock further facilitates the acceleration thanks to the sharp jump of physical parameters across the shock surface (Kong et al., 2019). We refer interested readers to Appendix C and references therein for more technical details.
@@
Figure 9: Comparison between the observed and modeled microwave images and spectra. (a) Observed multifrequency microwave images overlaid on the SDO/AIA 131 Å EUV image (identical to Figure 8(f)). (b) Sample microwave spectra derived from four selected regions marked in (a). (c) and (d) are similar to (a) and (b) but instead show the modeled multifrequency microwave images and the corresponding sample spectra.
@@
By combining the resulting distribution of energetic electrons from the particle model $dn_{e}/d\varepsilon(x,y,\varepsilon)$ and plasma properties including magnetic field $B(x,y)$, thermal plasma density $n_{\rm th}(x,y)$, and temperature $T(x,y)$ from the MHD model, we generate synthetic SDO/AIA EUV, RHESSI HXR, and EOVSA microwave images at different energy/frequency bands. Appropriate emission mechanisms are assumed for each emission type: optically thin line emissions for EUV, thin-target bremsstrahlung for HXRs, and nonthermal gyrosynchrotron emission for microwaves. The corresponding instrument response is also considered (see Appendix D for details). Figure 8(e) shows a composite of the resulting synthetic SDO/AIA 131 Å EUV image (background), RHESSI 25–60 keV HXR image (red contours), and multifrequency EOVSA microwave images (filled color contours). All of the images display a striking resemblance to the observations shown in Figure 8(f): the EUV 131 Å image displays a bright, closed arcade with a cusp-shaped top. A compact HXR source is located in the magnetic bottle region near the fast-mode termination shock above the flare arcade. Meanwhile, similar to the observations, the multifrequency microwave source resembles an arcade-like shape at the outer rim of the flare arcade. Figure 5 shows a side-by-side comparison between the observed and modeled microwave images at the same frequencies. Figures 8(e) and (f) display another comparison with the microwave images shown as overlaid filled contours colored from red to blue for increasing frequencies. One can see that the model images resemble the observed ones remarkably well. Moreover, as shown in Figure 9, not only does the modeled microwave source morphology achieve an excellent match with the observations, but the spatially resolved microwave spectra derived from different regions also yield a qualitative agreement, demonstrating the success of our approach. The microwave source’s appearance is very different from its HXR counterpart mainly because, unlike the X-ray bremsstrahlung, gyrosynchrotron radiation has a strong dependence on the local magnetic field.
@@
Intriguingly, unlike the HXR source that is highly localized at the ALT magnetic bottle, similar to the observations, the multifrequency microwave source extends well beyond the region toward the directions of the current sheet and the two footpoints. Such a large extension suggests that, while the majority of the energetic electrons are confined within the ALT HXR source near the Y point—in the model, it contains $\sim$50of the total $>$50 keV electrons (see red contour in Figure 8(d))—a small fraction of these electrons manage to escape and spread beyond the magnetic bottle region (albeit “diluted” to a much smaller density), giving rise to the extended microwave source with an arcade-like shape.
@@
We note that in the observations, another strong loop-top X-ray source is present, whose centroid is located at $y\approx 24$ Mm (Figure 8(f)). Spectral analysis results suggest that it is extremely hot (15–27 MK) and dense (up to $10^{12}$ cm-3). This source has been interpreted as the result of thermalization of previously accelerated nonthermal particles (Veronig & Brown, 2004), colliding chromospheric upflows (Reeves et al., 2007), or compression by slow shocks formed by sheared reconnection (Longcope & Guidoni, 2011). Reproducing this additional source in a model requires the inclusion of feedback between the thermal plasma and nonthermal particles, more accurate treatment of the chromospheric evaporation processes, and modeling of the reconnection processes in the third dimension, which is beyond the scope of our current work.
@@ 4 Discussion and Conclusion
We have demonstrated that our model, which includes both MHD and particle processes in a realistic flare geometry, can reproduce emission signatures that are well matched to multiwavelength observations. In the model, the energetic electrons are found to be primarily accelerated in the magnetic bottle region by the converging flows and facilitated by the fast-mode termination shock via the Fermi mechanism. The accelerated electrons are further trapped there due to pitch-angle scattering by turbulence. This is the first time that synthetic observables in both the thermal and nonthermal regimes have been generated from a self-consistent, macroscopic numerical model to compare with microwave, EUV, and X-ray imaging and spectroscopy observations. Such a remarkable agreement between the modeled and observed emissions suggests that this model of electron acceleration and transport is a viable approach during this period of interest.
The presence of the magnetic bottle structure in the ALT region of this event is well supported by its lower magnetic field strength derived from the microwave imaging spectroscopy data (Figure 6), as well as the observed abrupt change in the morphology and speed of the EUV plasma downflows (Figure 7). Direct evidence for fast-mode termination shocks is more elusive due to various challenges in identifying them observationally (see, e.g., discussions in Chen et al. 2019). However, recently, new and convincing evidence for their existence started to emerge. For this particular event, by using EUV imaging spectroscopy data recorded by Hinode/EIS, French et al. (2024) reported a sharp gradient in the Doppler velocity of hot ($\sim$18 MK) EUV downflows in the ALT region. The location and characteristics of the sharp velocity gradient agree very well with an MHD model that features a fast-mode termination shock at the same location. In addition, Cai et al. (2019) argued that the hot “supra-arcade fan” structure observed by SDO/AIA, Hinode/EIS, and the Interface Region Imaging Spectrograph was possibly caused by a termination shock, albeit the reported structure is not located above the main flaring arcade but in a loop system south of that.
@@
Figure 10: Physical picture suggested by our combined observation–modeling results. Energetic electrons are accelerated via the Fermi mechanism and trapped in the ALT magnetic bottle region due to turbulence. The reconnected and relaxed magnetic loops filled with hot plasma are observed as a bright EUV flare arcade with a hot loop-top HXR source. The highly concentrated energetic electrons near the Y point produce the ALT HXR source via bremsstrahlung. Meanwhile, a small fraction of energetic electrons escaping from the magnetic bottle region produces the observed multifrequency microwave source with a large extension, which has an arcade-like shape bestriding above the EUV flare arcade.
@@
The escaped nonthermal electrons from the ALT region gyrate in the magnetic field and produce microwave sources that form an arcade-like shape bestriding above the bright EUV flare arcade. The absence of an HXR footpoint source in our case suggests that it is either partially occulted by the solar limb at the time or the flux of nonthermal electrons reaching the footpoints is insufficient to produce observable HXR footpoint emission in the presence of a bright loop-top source. In other cases where HXR footpoint sources are co-observed, they are often found to be located at the outer edge of bright EUV flare arcades (e.g., Liu, 2013; Krucker & Battaglia, 2014). These observations nicely corroborate our scenario: they result from accelerated nonthermal electrons escaping from the ALT magnetic bottle region and reaching the footpoints along the freshly reconnected field lines (see, e.g., recent modeling results by Kong et al., 2022b).
As depicted in Figure 10, most of the accelerated electrons are confined near the Y point, producing the observed ALT HXR source. Meanwhile, the full extension of the observed microwave source traces 0.1%–1% of the peak $>$50 keV nonthermal electron density, or merely a millionth of the background plasma density! Our observations demonstrate the extreme sensitivity of microwave emission to even a tiny population of flare-accelerated nonthermal electrons. Such a sensitivity partially benefits from the Razin effect (Ginzburg & Syrovatskii, 1965). This effect strongly suppresses the microwave brightness in regions with relatively low magnetic field strength and high thermal plasma density, which are the exact features of the magnetic bottle region in our case. Strong Razin suppression occurs below a critical frequency of $\nu_{c}\approx 20n_{\rm th}/B$, or $\approx$5–6 GHz in the magnetic bottle region. This suppression effect is likely responsible for the apparent “gap” in the observed and modeled microwave sources at around $x=0$ Mm for frequencies $\lesssim$6 GHz (see Figure 5). Also, it effectively reduces the microwave brightness of the magnetic bottle region despite its high concentration of nonthermal electrons and, in turn, facilitates the detection of microwave sources arising from an extremely small nonthermal electron population outside the region.
The strong concentration of nonthermal electrons within the magnetic bottle region has important implications for electron acceleration and transport processes. First, it is highly likely that the nonthermal electrons are primarily accelerated in the magnetic bottle region itself rather than injected from elsewhere, such as high up in the reconnection current sheet or low in the loop legs. As shown in Figure 8(d), our joint observation–modeling results suggest that the nonthermal electron density rapidly drops to $<$1% of the peak value outside the ALT magnetic bottle region in all directions. Such a large density contrast implies that if the injection-and-trapping process is responsible, it must be a slow process. For instance, assuming all the electrons are injected from the upper current sheet, pure injection of 50 keV electrons without any loss takes $\tau^{\rm inj}\approx N_{e}^{\rm ALT}/F_{e}^{\rm inj}=n_{e}^{\rm ALT}L_{y}^{\rm ALT}A_{xz}^{\rm ALT}/(n_{e}^{\rm CS}v^{\rm inj}A_{xz}^{\rm CS})\approx(n_{e}^{\rm ALT}/n_{e}^{\rm CS})(A_{xz}^{\rm ALT}/A_{xz}^{\rm CS})L_{y}^{\rm ALT}/v^{\rm inj}$, where $n_{e}^{\rm ALT}/n_{e}^{\rm CS}>100$ is the ratio of the nonthermal electron density between the ALT and current sheet region, $A_{xz}^{\rm ALT}/A_{xz}^{\rm CS}>50$ is the expansion factor of the cross section between the magnetic bottle region (with a width of $L_{x}^{\rm ALT}\approx 10$ Mm) and the narrow current sheet (with a width of $L_{x}^{\rm CS}\approx 0.2$ Mm, the grid size of the numerical model), $L_{y}^{\rm ALT}$ is the vertical extension of the ALT region taken to be $\sim$5 Mm as suggested in the model, and $v^{\rm inj}$ is the injection speed of the electrons, taking the kinetic speed of 50 keV electrons $v_{e}^{50}\approx 0.41c$ as the upper limit. The estimated injection timescale to produce the ALT source is $\tau_{\rm inj}>200$ s, even without any loss. This timescale is much longer than the typical acceleration timescale inferred from HXR and radio emissions, which can display rapid fluctuations at second or even subsecond scales (e.g., Fletcher et al. 2011 and references therein), although we note that the acceleration in the flare gradual phase may be less variable and, as such, may have an inherently longer timescale. While we cannot entirely rule out such an injection-and-trapping scenario owing to the insufficient observational constraints for the precise dimensions of the ALT region (particularly in the LOS $z$-direction), we conclude that local acceleration and trapping in this region is a more likely scenario to account for the profound concentration of the nonthermal electrons in the ALT magnetic bottle region.
Finally, our results provide new insights into understanding the puzzling departure from equipartition between the upward-escaped and downward-retained energetic electron population reported by studies that combine remote-sensing and in situ observations, which have concluded consistently that only 0.1–1% of the flare-accelerated energetic electrons manage to escape to interplanetary space (Lin, 1974; Krucker et al., 2007; Wang et al., 2021, 2023a). Such a large imbalance not only poses a challenge for understanding the particle acceleration and transport processes but also has important implications for space weather in solar and extrasolar systems. In our scenario, this imbalance naturally occurs because the primary acceleration site is the magnetic bottle located below the X point, with the newly reconnected, cusp-shaped field lines in the large-scale current sheet acting as a nearly closed structure to limit the upward-directed electrons from escaping (Figure 10). Along with efficient trapping, as in our case, the energetic electron density reaching the X point can be only $<$1% of that at the core acceleration region, giving rise to an extremely small fraction of these energetic electrons that enter interplanetary space.
We note, however, that since our model does not involve an exhaustive search in the parameter space and does not include all possible particle energization/transport mechanisms, it is by no means exclusive. Other scenarios that result in an efficient energization and confinement of nonthermal electrons in the magnetic bottle region while allowing only a small fraction of them to escape may also be possible. Some candidates may include magnetic islands (Drake et al., 2006; Guidoni et al., 2022), collapsing traps (Somov & Kosugi, 1997), and intense shock heating (Masuda et al., 1994; Mann et al., 2024), yet rigorous data–model comparisons are required to further examine these models. Moreover, in order to achieve a more definitive understanding, extensive studies are needed for a large sample of flare events with different intensities and geometries. Last but not least, to make further progress, next-generation telescopes capable of performing radio and HXR imaging spectroscopy with orders-of-magnitude higher dynamic range and sensitivity, as well as more sophisticated models, are desired. Notable future telescope concepts include the Frequency Agile Solar Radiotelescope (Chen et al., 2023; Gary et al., 2023) in radio wavelengths and the Focusing Optics X-ray Solar Imager (Christe et al., 2016) or its variants (e.g., COMPLETE; Caspi et al. 2023) in X-rays.
This work is primarily funded by NASA Heliophysics Supporting Research grant 80NSSC20K1318 to NJIT, SAO, UMN, and LANL. B.C. and S.Y. received additional support from NSF grants AST-1735405 and AST-2108853 to NJIT. The authors are grateful to Dr. Harry Warren for providing the Hinode/EIS temperature and emission measure maps. They also thank Drs. Alexey Kuznetsov and Gregory Fleishman for making their fast gyrosynchrotron codes available. X.K. is supported by NSFC grants 42074203 and 11873036. C.S. is supported by NASA grants 80NSSC19K0853 and 80NSSC21K2044 and NSF AST 2108438 to SAO. X.L. acknowledges the support from NASA through grant 80NSSC21K1313, NSF grant AST-2107745, SAO through subcontract SV1-21012, and LANL through subcontract No. 622828. The Expanded Owens Valley Solar Array (EOVSA) was designed and built and is now operated by the New Jersey Institute of Technology (NJIT) as a community facility. EOVSA operations are supported by NSF grant AGS-2130832 and NASA grant 80NSSC20K0026 to NJIT. The authors are grateful to the RHESSI, SDO, SOHO/LASCO, and GOES-R/SUVI teams for making their data publicly available.
Appendix A: X-ray data analysis
Figure 11: RHESSI X-ray images for the time of interest on 2017 September 10. The time interval used for imaging is 16:10:00–16:11:10 UT. The energy ranges used for imaging are 10–16 keV, 16–25 keV, and 25–60 keV, shown in panels (a), (b), and (c), respectively. The solid and dashed lines denote the 50% and 10% contours. Panel (d) shows X-ray images of all three energy ranges overlaid on the SDO/AIA 131 Å (Fe XXI) EUV image at the closest time (shown in reversed gray scale).
X-ray spectral analysis is performed using the ospex software (Schwartz et al., 2002). We use the data from detector 3 (which had the lowest sensitivity at the time and hence was less affected by the pileup effect) to obtain the X-ray count flux spectrum between 16:10:00 UT and 16:11:08 UT. The resulting X-ray photon flux spectrum after applying the instrument response matrix is shown in Figure 4 as a black curve. Spectral modeling is performed in the 12–60 keV range using two isothermal functions and one nonthermal bremsstrahlung model. Also included in the spectral modeling is the pileup module. We performed spectral analysis assuming both the thick- and thin-target regimes. In both regimes, the thermal components, which dominate the spectrum below $\sim$30 keV, include a hot $\sim$15 MK loop-top source with a volume emission measure of $\approx 2\times 10^{51}\,\mathrm{cm}^{-3}$. A secondary superhot component with a temperature of 27–28 MK and a volume emission measure of $\approx 4\times 10^{49}\,\mathrm{cm}^{-3}$ is also present. Assuming a source volume $V\approx(10\,\mathrm{Mm})^{3}$ according to the size of the source shown in the X-ray image, the plasma density for the two thermal components becomes $\approx 1\times 10^{12}$ cm-3 and $2\times 10^{11}$ cm-3, respectively. Such a high density associated with the hot 15 MK component is sufficient to stop all nonthermal electrons up to 70 keV in the source (with a half-width of $\sim$5 Mm) through Coulomb collisions (Tandberg-Hanssen & Emslie, 1988).